# M4GT-Bench: Evaluation Benchmark for Black-Box Machine-Generated Text Detection Аннотация

### v1

@@ 5.3 Multi-way Detection
We try to answer three questions: (1) given a detector, which domain is the hardest to predict; (2) which generator is the most challenging one to distinguish; and (3) which detector performs the best on task 2. As results in Table 7, we see that Transformer-based black-box detectors demonstrate poor performance in this task, when facing with text from new domains not included in their training data. The detectors exhibit significant challenges in generalizing to unfamiliar content, reflecting a notable weakness in their ability to handle previously unseen domains. The domain of arXiv is the hardest for most detectors to identify if they are trained without arXiv examples. For most domains, davinci-003 is the most challenging generator to distinguish followed by ChatGPT. BLOOMz appears to be the easiest one, for all classifiers (see class-wise F1-scores in Appendix C). Even the worst detector NELA-SVM can identify BLOOMz with >90% F1-score. This may result from the fact that the generative distribution of BLOOMz is significantly different from other generators, while the distribution between davinci-003 and ChatGPT are more similar to each other, making it hard to make distinctions. RoBERTa is the best in this task followed by XLM-R. Based on the GLTR features, SVM one-vs-one mode is better than logistic regression one-vs-rest mode. Under the same SVM, GLTR features are more helpful than stylistic and NELA.
@@ 5.4 Boundary Identification
The results for task 3 are presented in Table 8. When training and testing within the same domain, it is much harder for boundary detection models to predict changing points for unseen generators, e.g., when training on generations from LLaMA-2 and testing on ChatGPT and vice versa. Testing on an unseen domain (mixture of generators) results in worse results, from MAE<5 (under the setting of All for both detectors) to MAE>21 and > 15 for Longformer and Deterta-v3 respectively, especially for Longformer trained with LLaMA-2 data (MAE>53). This demonstrates challenges on unseen domains and generators, consistent with findings in task 1 and 2.
@@ 6 Conclusion and Future Direction
In this work, we extend M4 into a new benchmark M4GT-Bench for MGT detection, providing the widest coverage of nine languages, six domains, and nine state-of-the-art generators, including GPT-4 and LLaMA-2 series. We formulate three tasks, with a novel formulation of detecting the changing point from human-written to machine-generated. We observe that human readers are unable to consistently differentiate between generators and operate merely at the level of making random guesses. For all tasks, we find that the detector usually suffers a severe penalty in their performance when they are challenged with unseen domains or generators. Some generators such as BLOOMz and domains such as arXiv appear harder than others. Overall, we see that Transformer-based detectors such as RoBERTa and XLM-R are usually the best on average. In future work, we plan to develop a demonstration system assisting human to detect MGT, especially in soem high-stake domains such as clinical, finantial and laws, and distinguish the authorship of different LLMs, promoting the protection of intelligence property. In addition to text, we would extend to other modalities detection including speech, image and videos. From the perspective of adversarial attack such as paraphrasing, it is worthwhile to explore the mixed texts with multiple changing points or vague boundaries.
@@
| RoBERTa | All | 96.35 | 96.21 | 96.23 | 96.25 | 96.34 | 94.43 | 96.38 | 96.74 | 97.12 | 93.30 | 99.31 |
| arXiv | 60.55 | 41.11 | 33.67 | 41.11 | 44.91 | 0.49 | 27.97 | 64.61 | 2.28 | 38.91 | 56.49 | |
| PeerRead | 67.54 | 65.47 | 61.55 | 64.77 | 58.71 | 0.47 | 66.45 | 53.79 | 62.60 | 94.25 | 94.60 | |
| 73.59 | 67.97 | 62.33 | 67.97 | 89.89 | 10.38 | 50.14 | 29.12 | 84.72 | 77.12 | 94.92 | ||
| WikiHow | 70.94 | 67.60 | 64.91 | 64.20 | 27.95 | 39.25 | 50.45 | 86.94 | 67.46 | 83.72 | 98.57 | |
| Wikipedia | 40.44 | 50.10 | 37.61 | 49.64 | 0.22 | 1.02 | 0.53 | 64.30 | 64.67 | 55.78 | 76.74 | |
| XLM-R | All | 87.69 | 86.26 | 86.03 | 86.36 | 91.41 | 80.63 | 87.48 | 90.30 | 81.60 | 72.03 | 98.75 |
| arXiv | 51.40 | 41.73 | 38.82 | 41.73 | 52.99 | 10.15 | 73.11 | 7.06 | 39.07 | 40.30 | 49.08 | |
| PeerRead | 52.20 | 51.66 | 44.84 | 50.38 | 17.19 | 0.53 | 66.21 | 32.68 | 38.51 | 62.64 | 96.13 | |
| 69.82 | 60.58 | 59.92 | 60.58 | 75.21 | 48.82 | 49.27 | 37.89 | 58.66 | 52.55 | 97.02 | ||
| WikiHow | 64.01 | 60.81 | 57.38 | 56.13 | 43.67 | 55.08 | 26.79 | 64.95 | 66.35 | 46.13 | 98.67 | |
| Wikipedia | 58.57 | 41.07 | 37.92 | 40.25 | 8.42 | 40.51 | 31.51 | 36.79 | 38.28 | 48.21 | 61.70 | |
| GLTR-LR | All | 42.9 | 46.11 | 43.07 | 45.84 | 66.98 | 13.15 | 37.06 | 30.17 | 50.30 | 33.83 | 70.00 |
| arXiv | 21.5 | 30.61 | 23.19 | 30.60 | 55.28 | 0.31 | 1.45 | 3.24 | 30.07 | 32.22 | 39.80 | |
| PeerRead | 43.50 | 46.24 | 42.52 | 46.37 | 58.77 | 1.05 | 36.18 | 50.92 | 43.31 | 39.51 | 67.90 | |
| 46.88 | 47.91 | 43.90 | 47.91 | 84.05 | 22.77 | 19.58 | 35.48 | 62.80 | 13.20 | 69.43 | ||
| WikiHow | 43.87 | 38.78 | 38.28 | 37.23 | 57.49 | 14.69 | 28.41 | 11.35 | 54.79 | 18.79 | 82.44 | |
| Wikipedia | 41.91 | 35.66 | 35.54 | 34.08 | 67.54 | 11.78 | 14.10 | 20.12 | 32.22 | 33.90 | 69.14 | |
| GLTR-SVM | All | 53.19 | 53.47 | 52.50 | 53.75 | 70.15 | 32.16 | 53.37 | 42.57 | 53.57 | 41.41 | 74.28 |
| arXiv | 27.88 | 35.40 | 29.25 | 35.40 | 61.14 | 0.06 | 42.95 | 0.00 | 28.55 | 33.7 | 38.33 | |
| PeerRead | 33.98 | 37.18 | 34.27 | 37.54 | 60.61 | 0.11 | 31.57 | 1.29 | 31.45 | 41.64 | 73.21 | |
| 46.68 | 47.84 | 44.73 | 47.84 | 83.34 | 14.95 | 43.89 | 18.88 | 56.88 | 15.58 | 79.62 | ||
| WikiHow | 43.88 | 42.36 | 40.52 | 37.84 | 66.13 | 22.54 | 1.15 | 23.16 | 64.09 | 23.65 | 82.93 | |
| Wikipedia | 43.62 | 37.96 | 35.69 | 37.18 | 53.55 | 17.18 | 40.28 | 43.48 | 32.03 | 36.89 | 26.45 | |
| Stylistic-SVM | All | 78.57 | 37.80 | 47.86 | 35.53 | 45.16 | 18.85 | 46.54 | 56.60 | 48.02 | 26.48 | 93.36 |
| arXiv | 46.57 | 7.67 | 11.55 | 7.44 | 6.39 | 0.00 | 2.03 | 2.99 | 3.45 | 9.87 | 56.10 | |
| PeerRead | 53.40 | 27.16 | 30.39 | 25.30 | 17.27 | 0.58 | 6.15 | 60.72 | 5.64 | 33.37 | 89.01 | |
| 61.96 | 28.23 | 33.07 | 26.53 | 50.18 | 15.22 | 10.78 | 26.11 | 29.06 | 12.92 | 87.18 | ||
| WikiHow | 54.48 | 34.89 | 39.94 | 26.49 | 43.20 | 26.8 | 29.77 | 38.6 | 25.26 | 23.34 | 92.60 | |
| Wikipedia | 51.40 | 23.56 | 29.54 | 17.78 | 33.78 | 8.07 | 1.01 | 38.78 | 33.45 | 34.5 | 57.18 | |
| NELA-SVM | All | 66.99 | 27.48 | 34.80 | 25.60 | 37.64 | 0.36 | 18.63 | 30.52 | 48.00 | 14.75 | 93.72 |
| arXiv | 53.46 | 12.97 | 19.11 | 11.71 | 13.64 | 0.07 | 12.88 | 10.73 | 19.00 | 10.81 | 66.62 | |
| PeerRead | 48.53 | 24.76 | 22.84 | 23.31 | 7.80 | 0.00 | 5.1 | 41.17 | 9.59 | 4.72 | 91.50 | |
| 45.94 | 28.09 | 31.45 | 24.20 | 49.00 | 0.46 | 8.69 | 8.32 | 48.28 | 13.96 | 91.43 | ||
| WikiHow | 46.87 | 24.81 | 28.54 | 21.16 | 48.68 | 0.20 | 6.96 | 13.88 | 16.80 | 16.65 | 96.59 | |
| Wikipedia | 47.81 | 22.62 | 27.36 | 19.79 | 31.19 | 0.07 | 6.88 | 5.79 | 29.46 | 24.95 | 93.19 | |
@@
Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. |
| Human |
@@
Distractions while driving could lead to deaths of other civilians. There are reasons why this topic became an actual law. The law was just passed last year that is you’re on your phone while driving, that is an immediate ticket. Due to your criminal record (if it’s bad) you can even do jail time if you feel you do not apply to this law. The cops of our cities are doing their job by stopping people who does not follow this law. The main reason why this show not be condoned is, because most of these situations a citizen could get in a car accident. Approximately 1.6 million citizens die due to being on their cell phone while operating a vehicle. Nearly 390,000 people get real serious injuries due to texting and driving. One out of four of every car accident in the United States are due to texting and driving. In my opinion I believe texting and driving is the main reason we have so many deaths every year. texting and driving needs to be/ stay illegal. That way people will not do it so often, which could decrease the death rate each year. Texting and driving could possibly be the reason why our death rates increase over time. Especially in this day and age, because almost everything now is technology based I believe that texting and driving is not safe for yourself or others, so technology should be the last thing you are on while behind and active vehicle. Texts, calls, emails all can wait until you are in a secure place (not behind the wheel) to respond, and also that way nobody could die or have an injury. |
| Prompt |
@@
Consider the following problem statement: Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. Write a well-structured essay of approximately 325 words expressing your clear opinion on the issue. Your essay should provide a thorough analysis of the problem, explore potential implications, and propose viable solutions or recommendations. Be sure to consider different perspectives surrounding the problem and support your opinions with relevant evidence, examples, or logical reasoning. Additionally, emphasize the importance of addressing the issue and the potential impact of implementing your proposed solutions. Essay: |
| Machine |
@@
Distractions while driving are a multi-faceted danger, with texting and driving sitting at the proverbial epicenter. The amalgamation of visual, manual, and cognitive distractions involved in this act makes it a leading hazard on the roads. This essay elucidates the imperative nature of legislation against texting and driving, buttressed by statistics and personal outlook, and explores the ramifications of non-compliance alongside potential resolutions. To initiate, consider the statistical landscape painting the grim reality. The National Highway Traffic Safety Administration (NHTSA) reports that in the United States alone, approximately 3,000 people lose their lives annually in crashes attributed to distracted driving. Texting is particularly pernicious; it takes the driver2̆019s eyes off the road for an average of five seconds, amounting to blind traversal of the length of a football field at 55 miles per hour. This nexus of inattention and control relinquishment is the breeding ground for tragedies. Moreover, the moral gravity of the situation cannot be overstated. When a driver engages in texting, they tacitly place their transient communication above the safety of fellow road users. This reflects a negligent societal value towards human life which by any humane standard is inexcusable. Hence, illegalizing such behavior manifests a collective ethos valuing life and responsibility. However, legal proscription must be complimented by earnest enforcement. Weak adherence to anti-texting laws, muddled by insubstantial sanctions, can subvert the intended deterrent effect. I assert that the consequence for distracted driving should encompass hefty fines, driving license suspensions, and, for repeat offenders or resulting in severe injury or death, significant custodial sentences. To mitigate the urge to reach for the phone, technological solutions like smartphone applications that disable texting while driving, and robust in-vehicle systems, need broader promotion. Educational campaigns fortifying public awareness about the consequences of distracted driving, seem pivotal to shifting attitudes, especially among impressionable young drivers. A symbiosis of legal strictures, technology, and education presents a tripartite solution to this problem. In conclusion, driving demands undivided attention, and texting fractures this focus. Legislating against such distractions is not only a proactive step towards reducing traffic-related fatalities and injuries, but it also reinforces a societal norm that prioritizes collective safety over momentary convenience. The significance of addressing this issue is underscored by the unbearable cost of human lives that could have been saved. In the grand tableau of road safety, cutting down the text during driving is not just a law; it is an ethical imperative that should carry the weight of universal endorsement. |
@@
The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS |
| Human |
@@
We discuss the results from the combined IRAC and MIPS c2d Spitzer Legacy observations of the Serpens star-forming region. In particular we present a set of criteria for isolating bona fide young stellar objects, YSO’s, from the extensive background contamination by extra-galactic objects. We then discuss the properties of the resulting high confidence set of YSO’s. We find 235 such objects in the 0.85 deg2̂ field that was covered with both IRAC and MIPS. An additional set of 51 lower confidence YSO’s outside this area is identified from the MIPS data combined with 2MASS photometry. We describe two sets of results, color-color diagrams to compare our observed source properties with those of theoretical models for star/disk/envelope systems and our own modeling of the subset of our objects that appear to be star+disks. These objects exhibit a very wide range of disk properties, from many that can be fit with actively accreting disks to some with both passive disks and even possibly debris disks. We find that the luminosity function of YSO’s in Serpens extends down to at least a few x .001 Lsun or lower for an assumed distance of 260 pc. The lower limit may be set by our inability to distinguish YSO’s from extra-galactic sources more than by the lack of YSO’s at very low luminosities. A spatial clustering analysis shows that the nominally less-evolved YSO’s are more highly clustered than the later stages and that the background extra-galactic population can be fit by the same two-point correlation function as seen in other extra-galactic studies. We also present a table of matches between several previous infrared and X-ray studies of the Serpens YSO population and our Spitzer data set. |
| Prompt |
@@
Write an abstract of an article from ArXiv given its title so that it is indistinguishable from a genuine human-written abstract. Write in the style that most abstracts from ArXiv are written. Title: The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS. Abstract: |
| Machine |
@@
This paper presents the results of the Spitzer c2d survey of the Serpens star-forming region, focusing on the young stellar object (YSO) population as observed with the Infrared Array Camera (IRAC) and the Multiband Imaging Photometer for Spitzer (MIPS). The survey identified a significant number of YSO candidates, providing a comprehensive view of star formation in this region. The data obtained from IRAC and MIPS observations were used to construct color-color and color-magnitude diagrams, which were then utilized to classify the YSOs into different evolutionary stages. The spatial distribution of these objects was also analyzed, revealing a correlation with the dense molecular gas in the region. The results contribute to our understanding of the star formation process in large, nearby, interstellar clouds and provide a valuable resource for future studies. |
@@
| Title |
{RLtext} الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع |
| Human |
{RLtext} تجددت المعارك اليوم الجمعة بمحيط العاصمة الليبية طرابلس، وأفادت التقارير |

### v2

@@ 5.3 Multi-way Detection
We try to answer three questions: (1) given a detector, which domain is the hardest to predict; (2) which generator is the most challenging one to distinguish; and (3) which detector performs the best on task 2. As results in Table 7, Transformer-based detectors show poor performance in this task when facing with text from new domains not included in their training data. Taking RoBERTa as an example, low accuracy is presented when testing on any unseen domains, such as arXiv and Wikipedia with accuracy of 36 and 52. Only when training over all domains and testing over all domains, accuracy reaches 97. The detectors exhibit significant challenges in generalizing to unfamiliar content, reflecting a notable weakness in their ability to handle previously unseen domains. The domain of arXiv is the hardest for most detectors to identify if they are trained without arXiv examples. For most domains, davinci-003 is the most challenging generator to distinguish, followed by ChatGPT. BLOOMz appears to be the easiest one, for all classifiers (see class-wise F1-scores in Appendix C). Even the worst detector NELA-SVM can identify BLOOMz with >90% F1-score. This may result from the fact that the generative distribution of BLOOMz is significantly different from other generators, while the distribution between davinci-003 and ChatGPT are more similar to each other, making it hard to make distinctions. RoBERTa is the best in this task followed by XLM-R. Based on the GLTR features, SVM one-vs-one mode is better than logistic regression one-vs-rest mode. Under the same SVM, GLTR features are more helpful than stylistic and NELA.
@@ 5.4 Boundary Identification
In Table 8, when training on generations from LLaMA-2 PeerRead and testing on ChatGPT PeerRead and vice versa, even within the same domain, predicted boundary shows large gap with the gold label. This reveals that it is much harder for boundary detection models to predict changing points for unseen generators. Testing on an unseen domain OUTFOX with mixture of generators including ChatGPT, GPT-4 and LLaMA-2 series, results in worse MAE, i.e., column all test. Taking the setting of All as an example, MAE changes from <5 (column PeerRead LLaMA-2-7B* and PeerRead ChatGPT) to MAE>21 and >15 (column all test) respectively for Longformer and Deterta-v3. Particularly for Longformer trained with LLaMA-2 data, MAE is greater than 53 when testing on all test, demonstrating challenges on unseen domains and generators, consistent with findings in task 1 and 2.
@@ 6 Conclusion and Future Direction
In this work, we extend M4 into a new benchmark M4GT-Bench for MGT detection, providing a wide coverage of nine languages, six domains, and nine state-of-the-art generators, including GPT-4 and LLaMA-2 series. We formulate three tasks, with a novel formulation of detecting the changing point from human-written to machine-generated. We observe that human readers are unable to consistently differentiate between generators and operate merely at the level of making random guesses. For all tasks, we find that the detector usually suffers a severe penalty in their performance when they are challenged with unseen domains or generators. Some generators such as BLOOMz and domains like arXiv appear harder than others. Overall, Transformer-based detectors such as RoBERTa and XLM-R are usually the best on average. In future work, we plan to develop a demonstration system assisting human to detect MGT, especially in some high-stake domains such as clinical, financial and legal domains, and distinguish the authorship of different LLMs, promoting the protection of intelligence property. In addition to text, we would extend to other modalities detection including speech, image and videos. From the perspective of adversarial attacks such as paraphrasing, it is worthwhile to explore mixed texts with multiple changing points.
Acknowledgements
We thank the anonymous reviewers for their insightful suggestions.
@@ Acknowledgements
- Dugan et al. (2020) Liam Dugan, Daphne Ippolito, Arun Kirubarajan, and Chris Callison-Burch. 2020. RoFT: A tool for evaluating human detection of machine-generated text. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 189–196, Online. Association for Computational Linguistics.
- Dugan et al. (2023) Liam Dugan, Daphne Ippolito, Arun Kirubarajan, Sherry Shi, and Chris Callison-Burch. 2023. Real or fake text?: Investigating human ability to detect boundaries between human-written and machine-generated text. In Thirty-Seventh AAAI Conference on Artificial Intelligence, AAAI 2023, Thirty-Fifth Conference on Innovative Applications of Artificial Intelligence, IAAI 2023, Thirteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2023, Washington, DC, USA, February 7-14, 2023, pages 12763–12771. AAAI Press.
@@
RoBERTa | All | 96.96 | 97.01 | 96.94 | 97.00 | 96.77 | 96.53 | 97.75 | 96.22 | 97.96 | 94.00 | 99.34 |
arXiv | 55.72 | 36.55 | 32.29 | 36.55 | 32.70 | 1.61 | 33.88 | 47.59 | 8.48 | 38.55 | 63.24 | |
PeerRead | 70.58 | 70.12 | 66.89 | 69.47 | 65.40 | 0.20 | 76.10 | 72.83 | 62.05 | 95.45 | 96.22 | |
77.21 | 74.49 | 71.66 | 74.49 | 84.47 | 75.44 | 44.94 | 35.51 | 81.84 | 82.41 | 96.97 | ||
WikiHow | 72.62 | 70.56 | 68.85 | 68.36 | 38.91 | 41.52 | 64.73 | 90.41 | 62.32 | 85.09 | 98.97 | |
Wikipedia | 46.37 | 52.22 | 39.37 | 51.91 | 0.51 | 0.96 | 1.77 | 65.31 | 62.54 | 59.07 | 85.46 | |
OUTFOX | 71.71 | 65.04 | 66.40 | 78.25 | 72.63 | 57.80 | 87.00 | 0.00 | 73.46 | 74.28 | 99.63 | |
XLM-R | All | 90.73 | 90.37 | 90.16 | 90.17 | 89.79 | 87.32 | 90.10 | 93.34 | 88.54 | 82.40 | 99.66 |
arXiv | 51.29 | 43.88 | 41.71 | 43.88 | 65.03 | 9.65 | 75.14 | 2.73 | 35.99 | 42.07 | 61.40 | |
PeerRead | 53.68 | 52.14 | 46.10 | 50.91 | 20.69 | 9.09 | 66.43 | 34.33 | 34.77 | 63.92 | 93.50 | |
69.73 | 58.72 | 57.34 | 58.71 | 72.02 | 59.47 | 41.10 | 27.22 | 51.73 | 52.96 | 96.88 | ||
WikiHow | 65.73 | 60.84 | 58.45 | 57.38 | 42.71 | 59.17 | 36.53 | 66.30 | 57.54 | 48.27 | 98.65 | |
Wikipedia | 60.04 | 42.55 | 38.80 | 41.95 | 5.50 | 45.40 | 38.07 | 29.29 | 40.25 | 50.44 | 62.63 | |
OUTFOX | 51.94 | 42.44 | 43.00 | 52.10 | 40.71 | 56.11 | 57.98 | 0.00 | 3.27 | 43.44 | 99.50 | |
GLTR-LR | All | 42.36 | 43.96 | 40.32 | 45.06 | 65.13 | 11.98 | 40.22 | 14.39 | 44.88 | 33.0 | 72.65 |
arXiv | 26.24 | 34.45 | 26.92 | 34.45 | 58.38 | 0.06 | 17.78 | 0.34 | 37.59 | 35.84 | 38.47 | |
PeerRead | 42.2 | 44.1 | 39.04 | 44.32 | 60.13 | 0.49 | 48.37 | 19.08 | 37.83 | 40.61 | 66.77 | |
45.54 | 46.28 | 41.7 | 46.28 | 85.06 | 24.08 | 22.44 | 19.77 | 55.74 | 14.6 | 70.23 | ||
WikiHow | 41.86 | 39.39 | 38.24 | 38.74 | 59.14 | 13.92 | 32.06 | 1.03 | 57.84 | 19.52 | 84.16 | |
Wikipedia | 41.62 | 36.95 | 34.62 | 35.18 | 67.62 | 8.15 | 16.62 | 0.7 | 33.41 | 35.59 | 80.25 | |
OUTFOX | 29.68 | 32.38 | 29.18 | 38.78 | 49.79 | 12.24 | 37.5 | 0.0 | 1.71 | 17.51 | 85.49 | |
GLTR-SVM | All | 52.81 | 48.42 | 47.24 | 50.39 | 69.35 | 21.39 | 49.47 | 31.21 | 46.84 | 36.04 | 76.35 |
arXiv | 22.57 | 32.29 | 25.73 | 32.28 | 61.34 | 0.0 | 26.36 | 0.0 | 27.79 | 28.63 | 35.99 | |
PeerRead | 34.1 | 39.1 | 34.81 | 39.4 | 60.64 | 0.15 | 41.71 | 0.17 | 30.74 | 36.98 | 73.26 | |
43.21 | 46.19 | 40.94 | 46.19 | 84.84 | 0.6 | 38.21 | 9.83 | 53.87 | 19.74 | 79.53 | ||
WikiHow | 44.13 | 39.01 | 38.27 | 36.18 | 64.59 | 19.73 | 13.6 | 5.61 | 58.81 | 24.26 | 81.26 | |
Wikipedia | 34.86 | 30.95 | 26.72 | 29.81 | 51.83 | 2.62 | 34.37 | 0.0 | 30.19 | 37.92 | 30.09 | |
OUTFOX | 26.93 | 28.29 | 26.45 | 28.58 | 51.25 | 27.8 | 2.84 | 0.0 | 0.54 | 15.71 | 86.99 | |
Stylistic-SVM | All | 78.95 | 37.10 | 47.31 | 35.26 | 43.60 | 24.63 | 53.28 | 51.35 | 38.46 | 25.58 | 94.28 |
arXiv | 44.89 | 8.52 | 12.71 | 8.27 | 0.65 | 0.00 | 18.02 | 0.72 | 2.31 | 9.28 | 58.02 | |
PeerRead | 50.72 | 21.96 | 25.43 | 20.44 | 11.23 | 0.50 | 7.78 | 28.56 | 3.57 | 37.50 | 88.85 | |
60.98 | 27.25 | 31.16 | 24.43 | 44.59 | 20.84 | 9.80 | 15.15 | 21.17 | 14.86 | 91.72 | ||
WikiHow | 56.23 | 34.57 | 37.04 | 26.71 | 46.19 | 28.11 | 10.82 | 31.44 | 27.23 | 22.93 | 92.57 | |
Wikipedia | 47.94 | 21.21 | 27.77 | 16.06 | 33.62 | 8.46 | 1.77 | 31.31 | 22.78 | 30.43 | 66.03 | |
OUTFOX | 48.28 | 27.46 | 32.50 | 26.80 | 21.91 | 42.69 | 42.64 | 0.00 | 0.06 | 21.32 | 98.89 | |
NELA-SVM | All | 64.50 | 23.91 | 30.54 | 34.11 | 0.18 | 20.58 | 19.97 | 28.86 | 15.26 | 94.82 | |
arXiv | 47.35 | 11.53 | 16.20 | 10.94 | 4.34 | 0.00 | 18.55 | 3.80 | 7.32 | 11.67 | 67.74 | |
PeerRead | 44.63 | 20.00 | 20.97 | 18.72 | 9.83 | 0.00 | 9.58 | 20.16 | 7.53 | 7.95 | 91.72 | |
42.77 | 24.27 | 27.87 | 20.72 | 43.73 | 0.20 | 7.84 | 2.62 | 31.48 | 17.52 | 91.67 | ||
WikiHow | 48.33 | 25.81 | 25.51 | 21.83 | 47.41 | 0.13 | 1.59 | 2.52 | 15.15 | 15.11 | 96.63 | |
Wikipedia | 46.38 | 20.74 | 25.06 | 18.76 | 27.19 | 0.07 | 7.92 | 3.09 | 18.81 | 25.04 | 93.34 | |
OUTFOX | 35.05 | 17.44 | 18.48 | 19.18 | 2.00 | 0.20 | 24.84 | 0.00 | 0.00 | 2.66 | 99.68 |
@@
Problem Statement | Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. |
@@
Human | Distractions while driving could lead to deaths of other civilians. There are reasons why this topic became an actual law. The law was just passed last year that is you’re on your phone while driving, that is an immediate ticket. Due to your criminal record (if it’s bad) you can even do jail time if you feel you do not apply to this law. The cops of our cities are doing their job by stopping people who does not follow this law. The main reason why this show not be condoned is, because most of these situations a citizen could get in a car accident. Approximately 1.6 million citizens die due to being on their cell phone while operating a vehicle. Nearly 390,000 people get real serious injuries due to texting and driving. One out of four of every car accident in the United States are due to texting and driving. In my opinion I believe texting and driving is the main reason we have so many deaths every year. texting and driving needs to be/ stay illegal. That way people will not do it so often, which could decrease the death rate each year. Texting and driving could possibly be the reason why our death rates increase over time. Especially in this day and age, because almost everything now is technology based I believe that texting and driving is not safe for yourself or others, so technology should be the last thing you are on while behind and active vehicle. Texts, calls, emails all can wait until you are in a secure place (not behind the wheel) to respond, and also that way nobody could die or have an injury. |
@@
Prompt | Consider the following problem statement: Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. Write a well-structured essay of approximately 325 words expressing your clear opinion on the issue. Your essay should provide a thorough analysis of the problem, explore potential implications, and propose viable solutions or recommendations. Be sure to consider different perspectives surrounding the problem and support your opinions with relevant evidence, examples, or logical reasoning. Additionally, emphasize the importance of addressing the issue and the potential impact of implementing your proposed solutions. Essay: |
@@
Machine | Distractions while driving are a multi-faceted danger, with texting and driving sitting at the proverbial epicenter. The amalgamation of visual, manual, and cognitive distractions involved in this act makes it a leading hazard on the roads. This essay elucidates the imperative nature of legislation against texting and driving, buttressed by statistics and personal outlook, and explores the ramifications of non-compliance alongside potential resolutions. To initiate, consider the statistical landscape painting the grim reality. The National Highway Traffic Safety Administration (NHTSA) reports that in the United States alone, approximately 3,000 people lose their lives annually in crashes attributed to distracted driving. Texting is particularly pernicious; it takes the driver2̆019s eyes off the road for an average of five seconds, amounting to blind traversal of the length of a football field at 55 miles per hour. This nexus of inattention and control relinquishment is the breeding ground for tragedies. Moreover, the moral gravity of the situation cannot be overstated. When a driver engages in texting, they tacitly place their transient communication above the safety of fellow road users. This reflects a negligent societal value towards human life which by any humane standard is inexcusable. Hence, illegalizing such behavior manifests a collective ethos valuing life and responsibility. However, legal proscription must be complimented by earnest enforcement. Weak adherence to anti-texting laws, muddled by insubstantial sanctions, can subvert the intended deterrent effect. I assert that the consequence for distracted driving should encompass hefty fines, driving license suspensions, and, for repeat offenders or resulting in severe injury or death, significant custodial sentences. To mitigate the urge to reach for the phone, technological solutions like smartphone applications that disable texting while driving, and robust in-vehicle systems, need broader promotion. Educational campaigns fortifying public awareness about the consequences of distracted driving, seem pivotal to shifting attitudes, especially among impressionable young drivers. A symbiosis of legal strictures, technology, and education presents a tripartite solution to this problem. In conclusion, driving demands undivided attention, and texting fractures this focus. Legislating against such distractions is not only a proactive step towards reducing traffic-related fatalities and injuries, but it also reinforces a societal norm that prioritizes collective safety over momentary convenience. The significance of addressing this issue is underscored by the unbearable cost of human lives that could have been saved. In the grand tableau of road safety, cutting down the text during driving is not just a law; it is an ethical imperative that should carry the weight of universal endorsement. |
@@
Title | The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS |
@@
Human | We discuss the results from the combined IRAC and MIPS c2d Spitzer Legacy observations of the Serpens star-forming region. In particular we present a set of criteria for isolating bona fide young stellar objects, YSO’s, from the extensive background contamination by extra-galactic objects. We then discuss the properties of the resulting high confidence set of YSO’s. We find 235 such objects in the 0.85 deg2̂ field that was covered with both IRAC and MIPS. An additional set of 51 lower confidence YSO’s outside this area is identified from the MIPS data combined with 2MASS photometry. We describe two sets of results, color-color diagrams to compare our observed source properties with those of theoretical models for star/disk/envelope systems and our own modeling of the subset of our objects that appear to be star+disks. These objects exhibit a very wide range of disk properties, from many that can be fit with actively accreting disks to some with both passive disks and even possibly debris disks. We find that the luminosity function of YSO’s in Serpens extends down to at least a few x .001 Lsun or lower for an assumed distance of 260 pc. The lower limit may be set by our inability to distinguish YSO’s from extra-galactic sources more than by the lack of YSO’s at very low luminosities. A spatial clustering analysis shows that the nominally less-evolved YSO’s are more highly clustered than the later stages and that the background extra-galactic population can be fit by the same two-point correlation function as seen in other extra-galactic studies. We also present a table of matches between several previous infrared and X-ray studies of the Serpens YSO population and our Spitzer data set. |
@@
Prompt | Write an abstract of an article from ArXiv given its title so that it is indistinguishable from a genuine human-written abstract. Write in the style that most abstracts from ArXiv are written. Title: The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS. Abstract: |
@@
Machine | This paper presents the results of the Spitzer c2d survey of the Serpens star-forming region, focusing on the young stellar object (YSO) population as observed with the Infrared Array Camera (IRAC) and the Multiband Imaging Photometer for Spitzer (MIPS). The survey identified a significant number of YSO candidates, providing a comprehensive view of star formation in this region. The data obtained from IRAC and MIPS observations were used to construct color-color and color-magnitude diagrams, which were then utilized to classify the YSOs into different evolutionary stages. The spatial distribution of these objects was also analyzed, revealing a correlation with the dense molecular gas in the region. The results contribute to our understanding of the star formation process in large, nearby, interstellar clouds and provide a valuable resource for future studies. |
@@
Title | {RLtext} الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع |
Human | {RLtext} تجددت المعارك اليوم الجمعة بمحيط العاصمة الليبية طرابلس، وأفادت التقارير |